# Deep Generative Classification of Blood Cell Morphology

### v1

@@ 3.3 General Training Setup
3.4 Datasets
We utilised multiple datasets, including publicly available ones and a custom dataset, to develop and evaluate our diffusion classifier for haematological cell classification. Our custom dataset, which we will release publicly as soon as we have obtained the necessary ethics approvals, consists of 559,808 single cell images from 2,904 blood smear slides obtained from Addenbrooke’s Hospital in Cambridge, with a labelled subset of 5,000 images across 10 classes. Importantly, all images were anonymised before we received them, ensuring patient privacy and ethical compliance. These images were created using CellaVision, a specialised imaging technology for cellular analysis. Notably, we included an artefact class in our dataset, addressing a critical challenge in clinical applications. Blood smear slides often contain artefacts that can be mistaken for cells by deep learning-based cell detection models. By explicitly modelling these artefacts, our dataset aims to enhance clinical applicability.
A distinctive feature of our dataset is the inclusion of labeller confidence scores, an attribute rarely found in existing datasets. This addition provides valuable information for various analyses beyond simple correlations. For instance, it enables in-depth comparisons between model and human expert performance, particularly in how certainty relates to accuracy. Researchers can investigate the patterns of confidence exhibited by models and human experts across different cases, and identify specific scenarios where models might demonstrate different certainty-accuracy relationships compared to human experts. Such analyses offer insights into the strengths and limitations of both AI models and human experts in haematological cell classification. For a more detailed analysis of how labellers’ confidence levels correlate with each other see the supplementary material, Figure 9. Table 3 summarises the datasets used in our study, highlighting the unique attributes of our custom dataset.
Our custom dataset’s labelling process involved six haematology experts with {34, 25, 15, 12, 6, 5} years of experience in blood microscopy, respectively. Each expert labelled 1,000 images, including 200 common images for inter-rater reliability assessment. Experts assigned one of ten classes (basophil, blast, lymphocyte, monocyte, neutrophil, erythroblast, eosinophil, immature granulocyte, platelet, or artefact) and indicated their confidence level (High, Moderate, Low, or No Confidence) for each prediction. For images labelled by multiple experts, we used a majority vote to determine the final label. In case of a tie, the label provided by the expert with the longest experience in blood microscopy was used.
@@ 3.4 Datasets
For this initial phase, we employed a batch size of 64 and trained on an A100-80GB GPU for 76,000 steps. We refer to the model resulting from this first phase as our base model. In the second phase, we used our base model to train our 3-fold cross-validation models for 45,000 steps each.
Table 3: Summary of the datasets used in our experiments.
Dataset # Images # Classes Labeller Confidence Artefacts
Raabin-WBC [62] 17,965 5
LISC [63] 257 5
PBC [22] 17,092 8
Bodzas [43] 16,027 9
Ours 5,000 $9+1$*
    9 regular classes of blood cells plus 1 artefact class.
3.5 Turing Test
To assess the quality and authenticity of our fine-tuned diffusion model’s synthetic blood cell images, we conducted a Turing test with expert haematologists. This evaluation was designed to determine whether the model could effectively capture the underlying distribution of blood cell images across various cell types, a capability that traditional discriminative models are not inherently required to possess. Additionally, we sought to assess the accuracy of the generated cell types.
@@ 3.5 Turing Test
We fine-tuned CytoDiffusion on the combined PBC and Bodzas datasets, with all images resized to $512\times 512$ pixels. From this combined dataset, we randomly selected 1,400 images for validation, with the remaining images allocated to the training set. The model was trained for 44,000 steps, employing a limited augmentation strategy that included only random vertical and horizontal flips.
@@
Following the training phase, we generated a set of 144 synthetic images spanning all the classes using 50 inference steps. Each image was conditioned on one of nine blood cell types: basophil, blast, lymphocyte, monocyte, neutrophil, erythroblast, eosinophil, immature granulocyte, and platelet. For the Bodzas dataset, we merged the neutrophil segment and neutrophil band classes into a single neutrophil class and created a blast class, which comprised both lymphoblasts and myeloblasts. To create a balanced dataset for the Turing test, we randomly selected an equivalent set of 144 real images from our validation dataset, resulting in a total pool of 288 images.
Ten haematology specialists from our research group, with {34, 28, 25, 15, 10, 9, 6, 5, 5, 0} years of experience in blood microscopy, participated in the Turing test. Participants were informed that half of the images presented to them would be real images from our dataset, while the other half would be synthetic images generated by our model. Each specialist was presented with the 288 images in a randomised order and asked to perform two tasks: (1) Distinguish between synthetically generated and real images and (2) classify each image into one of the nine designated blood cell types.
3.6 In-Domain Performance
To establish a baseline for our diffusion-based classifier, CytoDiffusion, we evaluated its performance on standard in-domain classification tasks using three datasets: our custom dataset, Raabin-WBC, and PBC. For the Raabin-WBC dataset, we utilised the predefined Train split (90% for training, 10% for validation) and tested on the predefined TestA dataset. For the PBC dataset, we employed an 80–10–10 split for train–validation–test.
For our custom dataset, we employed a cross-validation approach. We divided the dataset into four folds and fine-tuned four separate models, enabling us to test on the full dataset. For each fold used in training, we further split the data into 80% for training and 20% for validation and we trained each model for 60,000 iterations.
For the Raabin-WBC and PBC datasets, we trained our model for 72,000 steps. For comparison, we also trained and evaluated EfficientNetV2-M and ViT-B/16 models.
Given the significant class imbalances present in these datasets (Raabin-WBC TestA ranges from 89 to 2,660 images per class, our custom dataset from 103 to 1,425, and PBC from 1,199 to 3,314), we report both overall accuracy and balanced accuracy.
3.7 Uncertainty Measure
To assess the relationship between model performance and uncertainty, we employed psychometric function analysis. This approach allows us to quantify how closely our model’s uncertainty measure approximates that of an ideal observer, exhibiting purely aleatoric uncertainty.
We utilised the cross-validated models trained on our custom dataset, as described in Section 3.6. For quantifying the model’s uncertainty, we calculated the difference between the two classes with the smallest error, rescaled in the interval $[0,1]$. To establish a measure of labeller uncertainty, we mapped the confidence levels provided by our expert haematologists to numerical values: 1.0 for High Confidence, 2/3 for Moderate Confidence, 1/3 for Low Confidence, and 0.0 for No Confidence. For images assessed by multiple labellers, we averaged their uncertainty scores.
Psychometric functions describe the relation between the performance of an observer and a (typically scalar) property of the observed. The performance of interest is usually detection or classification, expressed as a function of signal strength on a monotonically increasing scale. Since an ideal model is as confident as the data allows, exhibiting purely aleatoric uncertainty, we can quantify the proximity of a model to that ideal by fitting a psychometric function with model confidence as the index of signal strength. A good measure of uncertainty should conform closely to an ideal observer, yielding a sigmoid curve rising from a chance guess rate, $\gamma$, where uncertainty is maximal, to a lapse rate $\lambda$, where uncertainty is minimal and any errors are not explicable by insufficient information.
@@ 3.7 Uncertainty Measure
where $S$ is a sigmoid, here chosen as the logistic, scaled in the interval [0,1]. We use Bayesian inference, implemented in psignifit (https://github.com/wichmann-lab/psignifit under MATLAB 2022b), to obtain parameter estimates, with priors informed by the psychophysical literature [42]. For the threshold, we use a uniform prior over the range of the data with a cosine fall off to 0 over half the range of the data on either side of the maximum and the minimum, expressing a belief that the threshold is located with equal probability within the sampled range and may be up to 50% of the spread of the data outside that range with decreasing probability. For the width, we use a uniform prior between two times the minimal distance of two tested stimulus levels and the range of the stimulus levels with cosine fall offs to 0 at the minimal difference of two stimulus levels and at 3 times the range of the tested stimulus. For $\gamma$, we fix the parameter at 1/10, reflecting the 10 possible classes, and for $\lambda$ we use a beta-distribution with parameters (1,10). We report the threshold and width estimated parameters, and their posterior distributions, citing 95% credibility intervals.
Functions were fitted to CytoDiffusion performance with its uncertainty, with data from the entire test set; to individual expert performance with consensus expert uncertainty, with data from 200 cases for which consensus uncertainty was available; and to individual expert performance with CytoDiffusion model uncertainty, again with data from 200 cases for which consensus uncertainty was available. To provide sufficient support for each evaluated performance level, uncertainties were discretized into 25 bins for the CytoDiffusion function (where the full test set was available), and 12 bins for the expert functions (where 200 cases were available).
3.8 Anomaly Detection
To evaluate our model’s capability in detecting anomalous cell types, we designed an experiment that simulates real-world scenarios where rare or previously unseen cell types might appear in clinical samples. This approach involved excluding specific abnormal cell classes during training and assessing the model’s ability to identify these classes during testing.
We utilised two datasets for this experiment: the Bodzas dataset and the PBC dataset. For the Bodzas dataset, we merged the neutrophil segment and neutrophil band classes into a single neutrophil class and excluded the blasts class, which comprised both lymphoblasts and myeloblasts (5,036 images in total). From the PBC dataset, we excluded the erythroblast class (1,513 images). The data was split into 70% training, 10% validation, and 20% test sets, with all abnormal cells (blasts and erythroblasts) moved to their respective test sets.
@@ 3.8 Anomaly Detection
We used the uncertainty described in Section 1.3. To visualise each model’s ability to distinguish between normal and abnormal cells, we generated KDE curves of the confidence differences for both groups. To quantify this ability, we calculated the AUC for each model and dataset.
3.9 Domain Shift
To evaluate our model’s robustness under domain shift conditions, we leveraged the Raabin-WBC dataset [62] and the LISC dataset [63], following methodologies outlined in previous studies [31, 30]. We utilised the predefined Train split for training (90%) and validation (10%), and the TestB split as one of our test sets. The TestB split was created using a different microscope and camera type compared to the training and validation set, introducing a domain shift. Additionally, we employed the LISC dataset as a second test set, which was created using different microscope and camera types, as well as a different staining method, further increasing the domain shift challenge. To maintain consistency with previous studies [31, 30] and accommodate the lower resolution of LISC images, we resized all images to $224\times 224$ pixels. We used a batch size of 32 and trained for 22,000 steps on an NVIDIA RTX A5000 GPU.
For comparison, we also fine-tuned and tested EfficientNetV2-M and ViT-B/16. To ensure reliable performance evaluation, we repeated the training and evaluation process five times for each of the models (our diffusion-based model, EfficientNetV2-M, and ViT-B/16).
3.10 Efficiency
To assess the efficiency of our model in learning from a small number of images per class, we designed an experiment using the PBC dataset [22]. We randomly selected 250 images per class for the training set, 836 images per class for the test set, and the remaining images for the validation set. From the training set, we randomly sampled subsets of 10, 20, 50, or 150 images per class to simulate low-data scenarios.
We trained CytoDiffusion on an NVIDIA RTX A5000 GPU. For subsets of 10, 20, 50, and 150 images per class, we trained for 20,000, 40,000, 100,000, and 240,000 steps respectively, saving checkpoints every 1000, 3000, 7000, and 20000 steps. For each subset, we selected the checkpoint with the highest validation accuracy for testing.
For comparison, we also trained and evaluated EfficientNetV2-M and ViT-B/16 models using the same data subsets. To account for variability in the subsampling and initialisations, we repeated each experiment five times. This involved resampling the specified number of images per class (10, 20, 50, and 150) and retraining all models for each repetition.
3.11 Explainability
One of the integral aspects of CytoDiffusion’s utility in clinical settings is its ability to provide explainable predictions. To achieve this, we employ a counterfactual heatmap approach, which elucidates what changes would be necessary for an image to be classified under a different specified class. This method is particularly beneficial for understanding model decisions in complex medical imaging tasks such as the classification of blood cell types.
The process begins by calculating the difference between the original noise tensor, $\boldsymbol{\epsilon}$, and the noise tensor predicted by the model for each class condition $c$. This difference is recorded for all iterations, and the mean error tensor for each condition $c$ is computed as
$\Delta_{c}=\frac{1}{N}\sum_{n=1}^{N}\left(\boldsymbol{\epsilon}_{n}-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t_{n}},t_{n},c)\right)\,,$ (10)
where $N$ is the number of iterations.
Subsequently, for each class condition $c$, we calculate the deviation from the condition with the minimum error, designated as $\Delta_{\hat{c}}$, where $\hat{c}$ is the predicted class. The adjusted tensor, $\delta_{c}$, is then
$\delta_{c}=\Delta_{c}-\Delta_{\hat{c}}\,.$ (11)
Finally, these adjusted tensors $\delta_{c}$ are decoded back to the pixel space using the VAE decoder to obtain the counterfactual heatmaps
$\boldsymbol{H}_{c}=\mathcal{D}(\delta_{c})\,,$ (12)
where $\mathcal{D}$ denotes the VAE decoder, and $\boldsymbol{H}_{c}$ represents the heatmap for condition $c$. These heatmaps visually represent modifications that would shift the image classification from the predicted class to the target class $c$, thereby providing a powerful tool for explaining and validating model predictions.
Consortia
The members of the BloodCounts! consortium are: Martijn Schut, Folkert Asselbergs, Sujoy Kar, Suthesh Sivapalaratnam, Sophie Williams, Mickey Koh, Yvonne Henskens, Bart de Wit, Umberto D’Alessandro, Bubacarr Bah, Ousman Secka, Parashkev Nachev, Rajeev Gupta, Sara Trompeter, Nancy Boeckx, Christine van Laer, Gordon A. Awandare, Kwabena Sarpong, Lucas Amenga-Etego, Mathie Leers, Mirelle Huijskens, Samuel McDermott, Willem H. Ouwehand, James Rudd, Carola-Bibiane Schönlieb, Nicholas Gleadall and Michael Roberts.
Appendix A Ablation Studies on Weighting Functions and Pruning Hyperparameters
@@
Our findings revealed that whilst the median error scaled exponentially with time (with $k=5.5$ in our case, which performed better than 6 or 7), the standard deviation of the error did not follow this pattern. Instead, it formed a curve similar to a negative quadratic function. Given that our primary concern is relative differences rather than absolute values, we focused on this aspect. Consequently, we fitted a fourth-degree polynomial to the median of the standard deviation of the squared errors using data from the efficiency experiment with 20 images per class and used the reciprocal of this as our weighting function. The coefficients of our polynomial function are
@@
To evaluate these weighting functions, we conducted ablation studies where we trained a model as described in Section 1.6 with 20 images per class. During inference, we used {1, 5, 10, 20, 50, 100, 250, 500, 1000} trials per class without any pruning for each weighting function. Table 4 presents our results, showing the accuracy after 1000 trials and the mean accuracy across all trials. Our custom weighting function performed best, with the logistic regression and the normalised $\mathcal{N}(0,1)$ approaches following closely behind.
@@
Appendix B Generated Images
The images presented in Figure 8 showcase the capability of our fine-tuned diffusion model to generate highly realistic synthetic blood cell images. Each image in the grid represents one blood cell class. These synthetic images were part of the dataset used in the Turing test described in the main text, where expert haematologists were unable to reliably distinguish them from real blood cell images.
@@ Appendix B Generated Images
Appendix C Labeller Confidence Analysis
@@
Appendix D High-Level Explanation of Model Prediction

### v2

@@ Appendix B Training and Inference Parameters and Procedure
Appendix C CytoData Labelling Procedure
CytoData’s labelling process involved six haematology experts with {34, 25, 15, 12, 6, 5} years of experience in blood microscopy, respectively. Each expert labelled 1,000 images, including 200 common images for inter-rater reliability assessment. Experts assigned one of ten classes (basophil, blast, lymphocyte, monocyte, neutrophil, erythroblast, eosinophil, immature granulocyte, platelet, or artefact) and indicated their confidence level (High, Moderate, Low, or No Confidence) for each prediction. For images labelled by multiple experts, we used a majority vote to determine the final label. In case of a tie, the label provided by the expert with the longest experience in blood microscopy was used.
The uncertainty label allows for in-depth comparisons between model and human expert performance, particularly in how certainty relates to accuracy. Researchers can investigate the patterns of confidence exhibited by models and human experts across different cases, and identify specific scenarios where models might demonstrate different certainty-accuracy relationships compared to human experts. Such analyses offer insights into the strengths and limitations of both AI models and human experts in haematological cell image classification. For a more detailed analysis of how labellers’ confidence levels correlate with each other see Supplementary Figure 9. Table 3 summarises the datasets used in our study, highlighting the unique attributes of CytoData.
@@ Appendix C CytoData Labelling Procedure
For this initial phase, we employed a batch size of 64 and trained on an A100-80GB GPU for 76,000 steps. In the second phase, we started our training from this initial model to train our 3-fold cross-validation models for 45,000 steps each.
Appendix D Authenticity Test Setup
@@
We fine-tuned the model on the combined PBC and Bodzas datasets, with all images resized to $512\times 512$ pixels. From this combined dataset, we randomly selected 1,400 images for validation, with the remaining images allocated to the training set. The model was trained for 44,000 steps, employing a limited augmentation strategy that included only random vertical and horizontal flips.
@@
Following the training phase, we generated a set of 144 synthetic images spanning all the classes using 50 inference steps and a guidance scale of 5, where the guidance scale controls the trade-off between image quality and adherence to the class condition [63]. Each image was conditioned on one of nine blood cell types: basophil, blast, lymphocyte, monocyte, neutrophil, erythroblast, eosinophil, immature granulocyte, and platelet. For the Bodzas dataset, we merged the neutrophil segment and neutrophil band classes into a single neutrophil class and created a blast class, which comprised both lymphoblasts and myeloblasts. To create a balanced dataset for the authenticity test, we randomly selected an equivalent set of 144 real images from our validation dataset, resulting in a total pool of 288 images.
Appendix E Psychometric Analysis Details
@@
where $S$ is a sigmoid, here chosen as the logistic, scaled in the interval [0,1]. We use Bayesian inference, implemented in psignifit (https://github.com/wichmann-lab/psignifit under MATLAB 2022b), to obtain parameter estimates, with priors informed by the psychophysical literature [45]. For the threshold, we use a uniform prior over the range of the data with a cosine fall off to 0 over half the range of the data on either side of the maximum and the minimum, expressing a belief that the threshold is located with equal probability within the sampled range and may be up to 50% of the spread of the data outside that range with decreasing probability. For the width, we use a uniform prior between two times the minimal distance of two tested stimulus levels and the range of the stimulus levels with cosine fall offs to 0 at the minimal difference of two stimulus levels and at 3 times the range of the tested stimulus.
Functions were fitted to CytoDiffusion performance with its uncertainty, with data from the entire test set; to individual expert performance with consensus expert uncertainty, with data from 200 cases for which consensus uncertainty was available; and to individual expert performance with model uncertainty, again with data from 200 cases for which consensus uncertainty was available. To provide sufficient support for each evaluated performance level, uncertainties were discretised into 25 bins for the CytoDiffusion function (where the full test set was available), and 12 bins for the expert functions (where 200 cases were available).
Appendix F Anomaly Detection Details
@@
For each image, we calculate the error for each class and normalise these errors by dividing by the mean error across all classes for that image. The confidence measure is then derived by computing the mean difference between the smallest error and all other errors, and subsequently subtracting the mean difference amongst all other errors. A larger score indicates higher confidence, as it suggests a greater distinction between the best-fitting class and the others. For the ViT, we used the logits (before the softmax) instead of the errors.
Appendix G Ablation Studies on Weighting Functions and Pruning Hyperparameters
@@
Our findings revealed that whilst the median error scaled exponentially with time (with $k=5.5$ in our case, which performed better than 6 or 7), the standard deviation of the error did not follow this pattern. Instead, it formed a curve similar to a negative quadratic function. Given that our primary concern is relative differences rather than absolute values, we focused on this aspect. Consequently, we fitted a fourth-degree polynomial to the median of the standard deviation of the squared errors using validation data from the efficiency experiment with 20 images per class and used the reciprocal of this as our weighting function. The coefficients of our polynomial function are
@@
To evaluate these weighting functions, we conducted ablation studies where we trained a model as described in the “Efficiency in Low-Data Regimes” section, with 20 images per class. During inference, we used {1, 5, 10, 20, 50, 100, 250, 500, 1000} trials per class without any pruning for each weighting function. Table 4 presents our results, showing the accuracy after 1000 trials and the mean accuracy across all trials. Our custom weighting function performed best, with the logistic regression and the normalised $\mathcal{N}(0,1)$ approaches following closely behind.
@@
Appendix H Generated Images
The images presented in Figure 8 showcase the capability of our fine-tuned diffusion model to generate highly realistic synthetic blood cell images. Each image in the grid represents one blood cell class. These synthetic images were part of the dataset used in the authenticity test described in the main text, where expert haematologists were unable to reliably distinguish them from real blood cell images.
@@ Appendix H Generated Images
Appendix I Labeller Confidence Analysis
@@
Appendix J High-Level Explanation of Model Prediction